# Cybersecurity threat detection based on a UEBA framework using Deep Autoencoders

### v1

@@ Abstract
User and Entity Behaviour Analytics (UEBA) is a broad branch of data analytics that attempts to build a normal behavioural profile in order to detect anomalous events. Among the techniques used to detect anomalies, Deep Autoencoders constitute one of the most promising deep learning models on UEBA tasks, allowing explainable detection of security incidents that could lead to the leak of personal data, hijacking of systems, or access to sensitive business information. In this study, we introduce the first implementation of an explainable UEBA-based anomaly detection framework that leverages Deep Autoencoders in combination with Doc2Vec to process both numerical and textual features. Additionally, based on the theoretical foundations of neural networks, we offer a novel proof demonstrating the equivalence of two widely used definitions for fully-connected neural networks. The experimental results demonstrate the proposed framework capability to detect real and synthetic anomalies effectively generated from real attack data, showing that the models provide not only correct identification of anomalies but also explainable results that enable the reconstruction of the possible origin of the anomaly. Our findings suggest that the proposed UEBA framework can be seamlessly integrated into enterprise environments, complementing existing security systems for explainable threat detection.
Keywords Anomaly detection $\cdot$ UEBA  $\cdot$ Autoencoders  $\cdot$ Deep learning  $\cdot$ Cybersecurity  $\cdot$ Cyber threat
@@ 1 Introduction
In the current digital era, cybersecurity and the reliability of both physical and logical systems have become of increased importance for industry and academia alike. The exponential growth of interconnected devices, the increasing volume of sensitive data, and the complexity of technological infrastructures highlight the need for robust algorithms that ensure security and resilience. Mathematical models and methods play a central role in this task by offering formal frameworks to identify cyberattacks, develop cryptographic protocols, simulate potential incidents under a wide range of scenarios, and design defence strategies against adversarial threats. As these challenges intensify, mathematical approaches are becoming more crucial to guarantee robust system performance and to mitigate future cyber risks.
@@
User and Entity Behaviour Analytics (UEBA) Shashanka et al. (2016a) is a powerful methodology for identifying cyber threats by creating models of normal behavioural patterns and detecting deviations that may indicate malicious or negligent activities. To profile the behaviour of an entity, these models allow the incorporation of multiple sources of information such as sensor readings, network traffic, system logs, security alerts, email information, and even geo-positioned or biometric data. UEBA uses advanced statistical learning techniques to model the behaviour of users, employees, and customers, as well as machines, such as servers, switches, and personal systems. By analysing anomalies in the behaviour of users and devices, UEBA can detect intrusions, impersonation attacks, or negligent users Maher (2017).
@@
In parallel, Explainable Artificial Intelligence (xAI) has emerged as a key enabler in cybersecurity scenarios by addressing the growing need for interpretability and trust in complex AI-driven security systems. Since modern cybersecurity solutions are relying on deep learning models, the inherent lack of transparency can prevent the analysts’ ability to understand and respond to the detected threats effectively. xAI techniques enable cybersecurity experts to interpret and understand why a system has flagged a specific event as anomalous behaviour, facilitating root-cause analysis and informed decision-making. Therefore, by integrating xAI techniques, UEBA-based cybersecurity tools not only improve technical performance but also enhance compliance with regulatory frameworks that demand transparency in AI-based decision-making, especially in sensitive areas such as finance.
@@
Several methods have been proposed in the literature related to the use of these techniques. In Shashanka et al. (2016b), UEBA models based on Mahalanobis distance and Singular Value Decomposition are implemented to identify anomalous behaviour in users accessing a server. Voris et al. Voris et al. (2019) use Gaussian mixture models for each computer, gathering file system, process launching, and network behaviour data, in addition to setting up a series of trap files to attract and identify attackers. They also apply UEBA to continuously identify the user in the system by monitoring their activity. Another example is found in Pusara and Brodley (2004), where decision trees are applied to mouse movement data to identify the user. Similarly, in Slipenchuk and Epishkina (2019), results of applying UEBA with mouse movement data, keyboard typing dynamics, and event sequences in the context of online banking operations are compared. In Meng et al. (2018), a similar approach based on the Radial Basis Function Network (RBFN) classifier with Particle Swarm Optimisation (PSO) is applied to verify user identity through touchscreen usage and other biometric data during web browsing. Moreover, the combination of static (logins, cookies, system type, etc.) and dynamic (mouse, keyboard, microphone, network usage, etc.) data to build user models was explored in Martin et al. (2021), where UEBA models are also used to guarantee user coherence when performing authentication with identity federations.
@@
UEBA can be considered a use case for anomaly detection (or outlier detection) with personalised models. Once the users are identified, each model has to detect data points that do not conform to the expected behaviour. As anomaly detection has grown in popularity, a wide array of methods and techniques has emerged. Among these, autoencoders Rumelhart et al. (1986) are a type of artificial neural network used for this purpose. Their goal is to learn (in an unsupervised way) a representation of the dataset by filtering out insignificant data or noise. Recent work in xAI for cybersecurity highlights the necessity of designing inherently interpretable (ante-hoc) models, prioritizing explainability principles from model conception through training Ortega-Fernandez et al. (2024). A key advantage of the use of autoencoders for cybersecurity is that they tend to be more explainable than other deep learning models Morales-Forero and Bassetto (2019); Gonzalez-Muniz et al. (2022). Moreover, in cases where the Well-Defined Anomaly Distribution (WDAD) assumption does not hold Görnitz (2019); Flynn et al. (2023), autoencoders can be trained on data assumed to be normal (even if slightly contaminated) Mauritz et al. (2021).
@@
Autoencoders have been used for anomaly detection since the work of Hawkins et al. Hawkins et al. (2002), and a general overview can be found in Wang et al. (2019). Examples include using Variational Autoencoders (VAEs) to construct anomaly scores Xiao et al. (2020) and convolutional autoencoders for video signal anomaly detection Ribeiro et al. (2018). In Sakurada and Yairi (2014) both incomplete and overcomplete autoencoders are applied in satellite data, while Wang et al. (2022) proposes a VAE coupled with a transformer architecture to account for dependencies in satellite data. In industrial anomaly detection, Zhou and Paffenroth (2017) uses a norm-regularised autoencoder, and Morales-Forero and Bassetto (2019) combines an autoencoder with an LSTM network. Also, Gonzalez-Muniz et al. (2022) utilises VAEs to classify anomalies in engineering systems. Finally, Ortega-Fernandez et al. (2023) showed that autoencoders outperform other algorithms in the detection of Denial of Service cyberattacks in industrial scenarios.
@@
Our work makes a twofold contribution. First, we provide novel theoretical results by proving the equivalence of two common definitions of fully-connected neural networks. Second, to the best of our knowledge, we present the first implementation of an explainable UEBA-based anomaly detection framework using autoencoders. Our methodology includes the use of text encoding models (Doc2Vec) alongside autoencoders to leverage both numerical and textual data, training with contaminated (unlabelled) data, and using model residuals for explainability.
@@
This paper is structured as follows. Section 2 details the proposed methodology, including theoretical results and a description of the methods employed for feature extraction, residual space analysis and the proposed architecture for UEBA-based anomaly detection. Section 3 presents the results of the application of the proposed methodology to a real use-case of cybersecurity in a financial institution. Finally, Section 4 outlines the main conclusions and future research directions.
@@ 2.1 Neural Networks
Neural networks constitute a large set of learning models that originate from the early work on the Rosenblatt perceptron Rosenblatt (1958). They approximate functions by interleaving affine transformations with non-linear activation functions. A feed-forward deep neural network use longer chains of concatenated affine and activation functions to improve the representation of the target function. This definition of a feed-forward deep neural can be formally expressed in the following manner.
@@
with each $\varphi_{j}\in\mathcal{F}$ for $j=1,\ldots,n^{(l)}$. Then, a feed-forward deep neural network with $d-1$ hidden layers is the function
@@
However, as noted in Mhaskar and Poggio (2019), this definition does not uniquely determine the network structure and makes it difficult to formalise concepts such as sparsity and convolutions. An alternative, more constructive description is based on a layered graph where each node (or neuron) implements a simple function, as formalized by Mhaskar and Poggio (2016); Cano-Córdoba et al. (2017). For brevity, we omit the definitions layered graph and neuron and they can be found in Cano-Córdoba et al. (2017).
@@
Neural networks of fixed depth or fixed width can approximate a wide range of functions modeling real-life processes Leshno et al. (1993); Kidger and Lyons (2020). The universality property is crucial for any application; hence, it’s significant to prove that both definitions are equivalent. Below, we provide a mathematical proof (Proposition 1) demonstrating this equivalence:
@@ 2.3 Doc2Vec
To process text-based variables (e.g., lists of executed processes), we use the Doc2Vec model Le and Mikolov (2014). Doc2Vec is a neural network-based embedding method that learns vector representations of documents in an unsupervised manner. Similar to Word2Vec Mikolov et al. (2013), it clusters similar texts in the vector space. In the same way that Word2Vec embeds related words like synonyms or topics close in the vector space, Doc2Vec also clusters similar texts together, for example by identifying the topic of the text or by finding similar words between texts. Two main algorithms exist for training Doc2Vec: Distributed Bag of Words (DBOW) and Distributed Memory (DM). In DBOW, the document vector is used to predict random word vectors from the document; in DM, both the document vector and word vectors are used to predict the next word in a sequence. The resulting document embeddings capture semantic similarities that are later used in our UEBA framework.
In this work, we apply Doc2Vec trained with DBOW to extract information from the lists of processes recovered from the activity logs. This way, we can capture when programs are run at the same time, and similarities in processes from their executable path.
2.4 t-distributed Stochastic Neighbor Embedding
The t-distributed Stochastic Neighbor Embedding (t-SNE) Van der Maaten and Hinton (2008) is a dimensionality reduction technique used primarily for visualizing high-dimensional data in two or three dimensions. It preserves local structures by mapping similar points from high-dimensional space to nearby points in the lower-dimensional embedding, while distant points remain separated. t-SNE works by first computing pairwise conditional probabilities based on distances between points in the high-dimensional space, and modelling the local similarities as Gaussian distributions. It then maps these points into a lower-dimensional space using a Student’s t-distribution to model the similarity of the embeddings. The algorithm iteratively adjusts embedding positions by minimizing the Kullback-Leibler divergence between these two distributions:
@@
where $P$ and $Q$ represent the joint probability distributions of pairwise similarities between data points in the high-dimensional and low-dimensional spaces, respectively. This lower dimensionality embedding allows us to identify patterns that are characterized by their local structure, such as clusters and anomalies in high dimensional datasets.
We use t-SNE to analyse the dataset and validate the autoencoder model’s behaviour. We apply t-SNE to the test data to analyse the presence of clusters indicative of different user groups and behaviours, and validate the feasibility of a UEBA-based approach. Moreover, we also use t-SNE to study the residuals of the reconstruction error of the model to verify that anomalous data points remain distinguishable within the residual space.

### v2

@@ Abstract
The increasing sophistication of cyberattacks, especially insider and process-related anomalies, poses a major challenge to enterprises, as traditional rule-based or shallow anomaly detection systems often fail to capture complex behavioural patterns. User and Entity Behaviour Analytics (UEBA) is a broad branch of data analytics that attempts to build a normal behavioural profile in order to detect anomalous events. Among the techniques used to detect anomalies, Deep Autoencoders constitute one of the most promising deep learning models on UEBA tasks, allowing explainable detection of security incidents that could lead to the leak of personal data, hijacking of systems, or access to sensitive business information. In this study, we introduce the first implementation of an explainable UEBA-based anomaly detection framework that leverages Deep Autoencoders in combination with Doc2Vec to process both numerical and textual features. Additionally, based on the theoretical foundations of neural networks, we offer a novel proof demonstrating the equivalence of two widely used definitions for fully-connected neural networks. The experimental results demonstrate the proposed framework’s capability to detect real and synthetic anomalies effectively generated from real attack data, showing that the models provide not only correct identification of anomalies but also explainable results that enable the reconstruction of the possible origin of the anomaly. Compared to existing UEBA and anomaly detection approaches, the novelty of our framework lies in combining explainable multimodal feature processing with formal mathematical guarantees. Our findings suggest that the proposed UEBA framework can be seamlessly integrated into enterprise environments.
Keywords Anomaly detection, User and Entity Behaviour Analytics, Autoencoders, Deep learning, Cybersecurity, Cyber Threat Detection
Mathematics Subject Classification: 68T07, 68M25
@@ 1 Introduction
In the current digital era, cybersecurity and the reliability of both physical and logical systems have become of increasing importance for industry and academia alike. The exponential growth of interconnected devices, the increasing volume of sensitive data, and the complexity of technological infrastructures highlight the need for robust algorithms that improve security and resilience. Mathematical models and methods play a central role in this task by offering formal frameworks to identify cyberattacks, develop cryptographic protocols, simulate potential incidents under a wide range of scenarios, and design defence strategies against adversarial threats. As these challenges intensify, mathematical approaches are becoming more crucial to guarantee robust system performance and to mitigate future cyber risks. Learning-based methods are increasingly employed to secure diverse environments. For example, in Internet of Things (IoT) systems, they are applied to detect attacks using heterogeneous sensor and network data [1]. In cyberphysical infrastructures, they have been explored to monitor and defend interconnected systems such as the energy-water nexus [2]. A related challenge arises in enterprise settings, where detecting anomalies in user activity —such as signs of cyberattacks, rogue insiders, or negligent behaviour— has become an essential task for modern security operations centres.
@@
User and Entity Behaviour Analytics (UEBA) [3] is a powerful methodology to identify cyber threats by creating models of normal behavioural patterns and detecting deviations that may indicate malicious or negligent activities. To profile the behaviour of an entity, these models allow the incorporation of multiple sources of information such as sensor readings, network traffic, system logs, security alerts, email information, and even geo-positioned or biometric data. UEBA uses advanced statistical learning techniques to model the behaviour of users, employees, and customers, as well as machines, such as servers, switches, and personal systems. By analysing anomalies in the behaviour of users and devices, UEBA can detect intrusions, impersonation attacks, or negligent users [4].
@@
In parallel, Explainable Artificial Intelligence (xAI) has emerged as a key enabler in cybersecurity scenarios by addressing the growing need for interpretability and trust in complex AI-driven security systems. Since modern cybersecurity solutions rely on deep learning models, its inherent lack of transparency can prevent the analysts’ ability to understand and respond to detected threats effectively. The integration of xAI techniques (such as SHAP, LIME or latent-space analysis) into cybersecurity operations enable cybersecurity experts to interpret and understand why a system has identified a specific event as anomalous behaviour, facilitating root-cause analysis and informed decision-making, prioritising genuine threats, and helping to identify false positives. Therefore, by integrating xAI techniques, UEBA-based cybersecurity tools not only improve technical performance, but also improve compliance with regulatory frameworks (such as the AI Act) that demand transparency in AI-based decision making, especially in sensitive areas such as finance.
@@
Several methods have been proposed in the literature related to the use of these techniques, which we summarise in Table 1. In [3], UEBA models based on Mahalanobis distance and Singular Value Decomposition (SVD) are implemented to identify anomalous behaviour in users accessing a server. Voris et al. [5] use Gaussian mixture models (GMM) for each computer, collecting data for the file system, process launch, and network behaviour, in addition to establishing a series of trap files to attract and identify attackers. They also apply UEBA to continuously identify the user in the system by monitoring their activity. Another example is found in [6], where decision trees are applied to mouse movement data to identify the user. Similarly, in [7], results of applying UEBA with mouse movement data, keyboard typing dynamics, and event sequences in the context of online banking operations are compared. In [8], a similar approach based on the Radial Basis Function Network (RBFN) classifier with Particle Swarm Optimisation (PSO) is applied to verify the user identity through touchscreen usage and other biometric data collected during web browsing. Moreover, the combination of static (logins, cookies, system type, etc.) and dynamic (mouse, keyboard, microphone, network usage, etc.) data to build user models was explored in [9], where UEBA models are also used to guarantee user coherence when performing authentication with identity federations.
@@
UEBA can be considered a use case for anomaly detection (or outlier detection) with personalised models. Once the users are identified, each model has to detect data points that do not conform to the expected behaviour. As anomaly detection has grown in popularity, a wide array of methods and techniques has emerged. Among these, autoencoders [10] are a type of artificial neural network used for this purpose. Their goal is to learn (in an unsupervised way) a representation of the dataset by filtering out insignificant data or noise. Recent work in xAI for cybersecurity highlights the necessity of designing inherently interpretable (ante-hoc) models, prioritizing explainability principles from model conception through training [11]. A key advantage of the use of autoencoders for cybersecurity is that they tend to be more explainable than other deep learning models [12, 13]. Moreover, in cases where the Well-Defined Anomaly Distribution (WDAD) assumption does not hold [14, 15], autoencoders can be trained on data assumed to be normal (even if slightly contaminated) [16].
@@
Autoencoders (AEs) have been used for anomaly detection since the work of Hawkins et al. [17], and a general overview can be found in [18]. Examples include using Variational Autoencoders (VAEs) to construct anomaly scores [19] and convolutional autoencoders for video signal anomaly detection [20]. In [21], both incomplete and overcomplete autoencoders are applied in satellite data, while [22] proposes a VAE coupled with a transformer architecture to account for dependencies in satellite data. In industrial anomaly detection, [23] uses a norm-regularised autoencoder, and [12] combines an autoencoder with an LSTM network. Also, [13] utilises VAEs to classify anomalies in engineering systems. Finally, [24] showed that autoencoders outperform other algorithms in the detection of Denial of Service cyberattacks in industrial scenarios.
Table 1: Comparison of related UEBA and anomaly-detection methods highlighting gaps addressed by our approach.
Method Domain Novelty / Differences w.r.t. this work
Shashanka et al. (2016) [3] UEBA Mahalanobis distance + SVD for user/device modelling; limited to numerical features without explainability; our work adds multimodal logs and interpretable residuals.
Voris et al. (2019) [5] UEBA GMM with decoys for active authentication; trap-based, not scalable to enterprise monitoring; our work scales to large heterogeneous logs.
Pusara & Brodley (2004) [6] Biometrics Mouse-movement re-authentication; narrow biometric focus; our work targets full enterprise behaviour.
Slipenchuk & Epishkina (2019) [7] UEBA Survey of UEBA statistical vs. ML methods; descriptive only, no unified DL framework; our work proposes a practical autoencoder pipeline.
Meng et al. (2018) [8] Biometrics Touch behaviour authentication with RBFN+PSO; mobile-focused, limited features; our work generalises to multimodal enterprise UEBA.
Martín et al. (2021) [9] UEBA UEBA for federated identity (OpenID Connect); restricted to identity scope; our work expands to enterprise-wide logs with textual features.
Morales-Forero & Bassetto (2019) [12] Industrial Semi-supervised AE for Industrial Control Systems (ICS) anomaly detection; domain-specific, no text integration; our work extends AE to enterprise logs with multimodal features.
González-Muñiz et al. (2022) [13] Industrial VAE with two-step residual classification for engineering systems; limited to industrial datasets; our work adapts VAE to enterprise UEBA with textual features and explainability.
Sakurada & Yairi (2014) [21] Satellite AE for anomaly detection in satellite telemetry; domain-limited; our work applies AE to enterprise UEBA with multimodal features.
Wang et al. (2022) [22] Satellite Likelihood-based AE for telemetry anomaly detection; strong in satellite domain, not generalisable; our work extends to enterprise logs with interpretability.
Zhou & Paffenroth (2017) [23] Industrial Robust deep AE for anomaly detection; designed for ICS data; our work adapts robust AE concepts to UEBA with text+numeric logs.
Ortega-Fernández et al. (2023) [24] Industrial Deep AE network intrusion detection system designed to detect distributed denial-of-service (DDoS) attacks in ICS; effective but network-traffic only; our work generalises AE to enterprise UEBA with multimodal logs and interpretability.
The motivation of this work lies in two key challenges of anomaly detection in enterprise environments. First, labelled attack data are rarely available, and the heterogeneity between infrastructures and behaviours limits the reuse of public datasets, which makes unsupervised approaches especially valuable. Second, most previous contributions focus only on numerical features, while enterprise logs also contain textual elements that are often ignored. In this work, we focus on executable paths as textual features, while noting that other sources, such as web addresses or email content, could also be explored in future research. In response, we propose a UEBA framework that integrates deep autoencoders with text embeddings, grounded on solid mathematical foundations.
@@
Our work makes a twofold contribution. First, we provide novel theoretical results by proving the equivalence of two common definitions of fully-connected neural networks. Second, to the best of our knowledge, we present the first implementation of an explainable UEBA-based anomaly detection framework using autoencoders. Our methodology includes the use of text encoding models (Doc2Vec) alongside autoencoders to leverage both numerical and textual data, training unlabelled (possibly contaminated) data, and using model residuals for explainability.
@@
This paper is structured as follows. Section 2 details the proposed methodology, including theoretical results and a description of the methods employed for feature extraction, residual space analysis, and the proposed architecture for UEBA-based anomaly detection. Section 3 presents the results of the application of the proposed methodology to a real use-case of cybersecurity in a financial institution. Finally, Section 4 outlines the main conclusions and future research directions.
@@ 2.1 Neural Networks
Neural networks constitute a large set of learning models that originate from the early work on the Rosenblatt perceptron [25]. They approximate functions by interleaving affine transformations with non-linear activation functions. A feed-forward deep neural network uses longer chains of concatenated affine and activation functions to improve the representation of the target function. This definition of a feed-forward deep neural network can be formally expressed in the following manner.
@@
with each $\varphi_{j}\in\mathcal{F}$ for $j=1,\ldots,n^{(l)}$. Then, a feed-forward deep neural network with d1d-1 hidden layers is the function
@@
However, as noted in [26], this definition does not uniquely determine the network structure and makes it difficult to formalise concepts such as sparsity and convolutions. An alternative, more constructive description is based on a layered graph where each node (or neuron) implements a simple function, as formalised by [27, 28]. For brevity, we omit the definitions of layered graph and neuron, and they can be found in [28].
@@
Neural networks of fixed depth or fixed width can approximate a wide range of functions modelling real-life processes [29, 30]. The universality property is crucial for any application; hence, it is significant to prove that both definitions are equivalent. Below, we provide a mathematical proof (Proposition 1) demonstrating this equivalence:
@@ 2.2 Autoencoders
It should be highlighted that Proposition 1 establishes that the two common formalisations of feed-forward neural networks (functional composition vs. layered graphs) are equivalent. Since our encoder and decoder are implemented as fully connected networks, this equivalence guarantees that the autoencoder used in our UEBA pipeline is a well-defined object in either formalism, with no loss of generality when moving between them. This allows us to apply the classical universal approximation results for feed-forward networks [29, 30], providing a rigorous foundation for our modelling choice: the autoencoder has sufficient expressivity to capture the dominant patterns of normal behaviour, while its residuals define a mathematically coherent and explainable anomaly score used throughout the framework.
@@ 2.3 Doc2Vec
To process text-based variables (e.g., lists of executed processes), we use the Doc2Vec model [31]. Doc2Vec is a neural network-based embedding method that learns vector representations of documents in an unsupervised manner. Similar to Word2Vec [32], it clusters similar texts in the vector space. In the same way that Word2Vec embeds related words like synonyms or topics close in the vector space, Doc2Vec also clusters similar texts together, for example, by identifying the topic of the text or by finding similar words between texts. Two main algorithms exist for training Doc2Vec: Distributed Bag of Words (DBOW) and Distributed Memory (DM). In DBOW, the document vector is used to predict random word vectors from the document; in DM, both the document vector and word vectors are used to predict the next word in a sequence. The resulting document embeddings capture semantic similarities that are later used in our UEBA framework by applying the Doc2Vec trained with DBOW to extract information from the lists of processes recovered from the activity logs. We specifically adopt Doc2Vec with the DBOW algorithm because executable paths behave more like structured categorical tokens than natural language, so co-occurrence at the window level captures useful behavioural similarity (e.g., programs launched together) without requiring heavy language models.
2.4 t-distributed Stochastic Neighbour Embedding
The t-distributed Stochastic Neighbour Embedding (t-SNE) [33] is a dimensionality reduction technique used primarily for visualising high-dimensional data in two or three dimensions. It preserves local structures by mapping similar points from high-dimensional space to nearby points in the lower-dimensional embedding, while distant points remain separated. t-SNE works by first computing pairwise conditional probabilities based on distances between points in the high-dimensional space, and modelling the local similarities as Gaussian distributions. It then maps these points into a lower-dimensional space using a Student’s t-distribution to model the similarity of the embeddings. The algorithm iteratively adjusts embedding positions by minimising the Kullback-Leibler divergence between these two distributions:
@@
where $P$ and $Q$ represent the joint probability distributions of pairwise similarities between data points in the high-dimensional and low-dimensional spaces, respectively. This lower dimensionality embedding allows us to identify patterns that are characterised by their local structure, such as clusters and anomalies in high dimensional datasets.
We use t-SNE to analyse the dataset and validate the autoencoder model’s behaviour. We apply t-SNE to the test data to analyse the presence of clusters indicative of different user groups and behaviours, and validate the feasibility of a UEBA-based approach. Moreover, we also use t-SNE to study the residuals of the reconstruction error of the model, verifying that anomalous data points remain distinguishable within the residual space.